# A Systematic Literature Review of Spatio-Temporal Graph Neural Network Models for Time Series Forecasting and Classification

### v2

@@ 5.7.5 Results
The most common error metrics used in the evaluation of the models in the “Generic” group are, in order, the mean absolute error (MAE), correlation coefficient (CORR), root mean squared error (RMSE), root relative squared error (RRSE), mean absolute percentage error (MAPE), and mean squared error (MSE). Differently from other metrics, where lower values indicate higher accuracy, the correlation coefficient (CORR) quantifies the strength of the the linear relationship between predicted and actual values, with higher values indicating a better performance. This subsection about results compares the results obtained for the Electricity consumption, Exchange rate, Solar energy and Traffic datasets. The results of the proposed models and benchmarks on METR-LA and PEMS datasets were already reported with a in Subsec. 5.5.5.
@@
DPGNN [art_405] 0.0720 0.9527 0.0810 0.9422 0.0893 0.9357 0.0935 0.9291
DSTGN [art_73] 0.0713 0.9518 0.0821 0.9424 0.0887 0.9357 - -
DSTIGNN [art_74] 0.0733 0.9515 0.0820 0.9412 0.0899 0.9338 0.0954 0.9289
ESG 0.0718 0.9494 0.0844 0.9372 0.0898 0.9321 0.0962 0.9279
@@
MAGNN [art_80] 0.0745 0.9476 0.0876 0.9323 0.0908 0.9282 0.0963 0.9217
MDST-GNN [art_90] 0.0738 0.9454 0.0833 0.9346 0.0884 0.9264 0.0922 0.9222
MTGNN 0.0745 0.9474 0.0878 0.9316 0.0916 0.9278 0.0953 0.9234
MTGODE [art_75] 0.0736 0.9430 0.0809 0.9340 0.0891 0.9279 - -
@@
The accuracy of the various models varies depending on the studied forecasting horizon but, overall, one can notice that the GNN models perform better than the others, with the static and dynamic graph learning network (SDGL) [art_78] being the most accurate. The SDGL model, proposed by Li et al. in 2023, first employs a static and dynamic graph learning module that incorporates graph regularization to control the smoothness, connectivity, and sparsity of the learned graph. The graph input then passes through gated temporal convolutional networks and graph convolution modules, which are equipped with residual and skip connections that deliver the information to the output layer.
@@
DPGNN [art_405] 0.0184 0.9797 0.0246 0.9715 0.0346 0.9564 0.0456 0.9386
DSTGN [art_73] 0.0179 0.9782 0.0250 0.9715 0.0346 0.9562 - -
DSTIGNN [art_74] 0.0173 0.9818 0.0244 0.9723 0.0333 0.9571 0.0430 0.9407
ESG 0.0181 0.9792 0.0246 0.9717 0.0345 0.9564 0.0468 0.9392
@@
MAGNN [art_80] 0.0183 0.9778 0.0246 0.9712 0.0343 0.9557 0.0474 0.9339
MDST-GNN [art_90] 0.0172 0.9811 0.0245 0.9727 0.0337 0.9578 0.0431 0.9392
MTGNN 0.0194 0.9786 0.0259 0.9708 0.0349 0.9551 0.0456 0.9372
@@
Also in this case, GNN models are the best. The so-called discrete graph structure learning for time series model (GTS), proposed in 2021 by Shang et al. in Ref. [GTS] and used as a benchmark in Ref. [art_78], demonstrates itself to have a superior performance. It is therefore a suitable benchmark for comparison. The GTS model was introduced by Shang et al. in 2021 for traffic and sensor network forecasting. It includes a probabilistic graph learning module that optimizes the mean performance over the graph distribution, and a recurrent graph convolutional module to generate the forecasts.
@@
DPGNN [art_405] 0.1794 0.9850 0.2364 0.9724 0.3003 0.9543 0.4091 0.9117
DSTGN [art_73] 0.1787 0.9867 0.2358 0.9721 0.3079 0.9513 - -
DSTIGNN [art_74] 0.1684 0.9869 0.2165 0.9773 0.2863 0.9584 0.4064 0.9118
ESG 0.1708 0.9865 0.2278 0.9743 0.3073 0.9519 0.4101 0.9100
@@
MAGNN [art_80] 0.1771 0.9853 0.2361 0.9724 0.3015 0.9539 0.4108 0.9097
MDST-GNN [art_90] 0.1764 0.9855 0.2321 0.9735 0.3082 0.9519 0.4119 0.9103
MTGNN 0.1778 0.9852 0.2348 0.9726 0.3109 0.9509 0.4270 0.9031
MTGODE [art_75] 0.1693 0.9868 0.2171 0.9771 0.2901 0.9577 - -
@@
The results indicate that the dynamic spatiotemporal interactive graph neural network (DSTIGNN) model [art_74] and the static and dynamic graph learning network (SDGL) model [art_78] are the most accurate on the solar energy dataset, for shorter and longer time horizons, respectively. The DSTIGNN model, introduced by Gao et al. in 2021, employs a hierarchical structure consisting of a spatio-temporal interactive learning module, a dynamic graph inference module, multiple spatio-temporal layers, and an output module. The spatio-temporal interactive learning module splits the time series into two subsequences to capture both temporal and spatial correlations. Then, the graph inference module learns a dynamic adjacency matrix for the graph. Once the graph structure is determined, a stack of graph convolutional layers, enhanced with residual connections, along with a convolutional output module, generate the final forecast. The description of the SDGL instead has already been provided in this subsection, in the discussion of the results on the Electricity consumption dataset.
@@
MAGNN [art_80] 0.4097 0.8992 0.4555 0.8753 0.4423 0.8815 0.4434 0.8813
MDST-GNN [art_90] 0.4162 0.8958 0.4461 0.8803 0.4377 0.8841 0.4452 0.8792
MTGNN 0.4162 0.8963 0.4754 0.8667 0.4461 0.8794 0.4535 0.8810
MTGODE [art_75] 0.4127 0.900 0.4259 0.8945 0.4329 0.8899 - -
@@
The accuracy of the models varies by horizon and metric. However, the best one in terms of RRSE over all horizons is the temporal decomposition enhanced graph neural network for multivariate time series forecasting (TDG4MSF) [art_81], proposed by Miao et al. in 2021. The TDG4MSF model is composed of four components: temporal decomposition enhanced representation learning for extracting short and long term periodic patterns, graph structure learning, gated graph neural network depending on linear transformation for representation learning, and MLP-based output layer.
@@ 5.8 Other topics
This final subsection includes 29 papers that fall outside the previously discussed groups. The case studies include cellular traffic prediction, recommendation systems for user preferences, and other forecasting or classification problems. Given the heterogeneity of the papers, only a general overview is provided here, and no comparisons of the approaches or discussions of benchmarks are presented.
@@ 5.8.1 Overview
The interest in GNNs has grown so extensively that the selected themes are actually not sufficient to cover all relevant application domains. However, since only a few papers would belong to each potential additional theme, they are grouped under the label “Other Topics”. A frequently discussed topic in this “Other topics” category is cellular traffic prediction [art_138, art_144, art_148, art_468, art_457], in terms of SMS messages, calls, internet connections, phone’s traffic statistics or locations, which are essential for the cellular network resource management system. Another topic included is modulation classification [art_12, art_13, art_303, art_459, art_463], which is the process of determining the modulation used at the transmitter based on observations of the received signal [modulation_classification_definition]. Two further areas of investigation to be added are the study of user preferences [art_147, art_151] to enable personalized experiences and recommendations, and anomalous IoT intrusion detection [4, art_460].
Other applications discussed in the papers of this subsection are encrypted traffic classification [art_118], urban spatio-temporal event prediction [art_136], prediction of the quality of operational processes [art_137], seismic intensity forecasting [art_139] (with source code available at https://github.com/StefanBloemheuvel/GCNTimeseriesRegression), prediction of spatio-temporal dynamics with complex structures [art_140], tailing dam monitoring [art_141], forecasting of educational video engagement [art_142], subsurface production forecasting [art_143], time series prediction for data centers maintenance [art_145], hydraulic runoff [art_146], water demand forecasting [art_149], human activity recognition [art_159] (with source code available at https://github.com/riktimmondal/HAR-Sensor ), crime prediction [art_465] (with source code available at https://github.com/ZJUDataIntelligence/HDM-GNN), polypropylene production data [art_455], and tungsten flotation process [art_458] (with source code available at https://github.com/ywmcsu/SASGNN).
@@ 5.8.2 Datasets
The two datasets for modulation classification used in Refs. [art_12, art_13, art_303, art_463] are RML2016.10a (https://www.kaggle.com/datasets/raindrops12/rml201610a, source paper: http://dx.doi.org/10.1007/978-3-319-44188-7_16) and RML2016.10b (https://www.kaggle.com/datasets/marwanabudeeb/rml201610b, source paper: https://pubs.gnuradio.org/index.php/grcon/article/view/11). They include 220,000 and 1.2 million samples respectively, in many modulation types. Each modulation type has 20 levels of signal-to-noise ratios at 2 dB intervals from -20 dB to 18 dB.
As for the cellular traffic prediction dataset used in Refs. [art_138, art_144, art_468], it was proposed by Telecom Italia and MIT Media Lab during the “Telecom Italia Big Data Challenge”, and it is available at https://dataverse.harvard.edu/dataverse/bigdatachallenge. The investigated geographical area is divided in a grid, and each cell records the number of SMS messages, calls, and wireless network traffic data in a 10 minutes interval.
Finally, the Bot-IoT dataset used in Refs. [4, art_460] and available at https://www.unsw.adfa.edu.au/unsw-canberra-cyber/cybersecurity/ADFA-NB15-Datasets/bot_iot.php, was created by designing a realistic network environment in the Cyber Range Lab of UNSW Canberra, and incorporates a combination of normal and botnet traffic.
@@ 6 Discussion
As stated in the Introduction, the aim of this SLR is to provide a comprehensive and detailed overview of spatio-temporal GNN models for time series classification and forecasting in different fields, and to assess their performance. Existing literature was collected and synthesized to provide an overview of datasets, models, and tables of results that will hopefully assist researchers in their future work. In the Introduction two sets of research questions were stated: a generic set and a specific set. This section will collect the answers to those research questions.
@@
SQ2 (Graph construction). Most of the selected papers focus on pre-defined graph structures (when available), with the objective of extracting the maximum amount of information and enhancing the interpretability of the model. However, there is a recent growing interest in models that learn the graph structure and the edge weights themselves. This trend is expected to become more prevalent in the future.
@@
SQ3 (Taxonomy). As observed in Ref. [survey_spatiotemporal], there are two main approaches for the design of spatio-temporal GNNs: one that treats the spatial and temporal components in separate modules, and another that integrates and processes them together. The analysis of the collected papers reveals that the most common approach in literature is that of separate modules. Specifically, researchers very often address the spatial and temporal aspects of the problem independently from each other, and focus on each module separately. Hence, the proposed taxonomy of GNN models takes its main perspective from the point of view of the spatial component. The review indicates that the convolutional and the attentional approaches are the prevalent ones. Approximately 62% of the models are purely convolutional, 21% are purely attentional, and 9% are hybrid convolutional-attentional. With regard to the temporal component, the recurrent structures of GRUs and the attention mechanism are widely used.
@@
SQ4 (Benchmark models). As for the benchmark models, there are many options available, and their choice depends on the specific application. In thematic groups such as “Energy”, “Finance”, “Health”, and “Predictive monitoring”, the focus tends to be on simpler statistical and non-GNN machine learning benchmarks, with only a limited use of GNN benchmarks. In the “Generic” group, many recent machine learning benchmark models based on the Transformer architecture have emerged. Notably, in “Mobility”, “Environment”, and “Generic” groups (which are also the most investigated fields), there are many reference GNN benchmark models. Among the most prevalent benchmarks there are: the graph convolutional neural network with long-short term memory (GC-LSTM) [GC-LSTM], spatio-temporal graph convolutional network (STGCN) [STGCN], attention-based spatial-temporal graph convolutional network (ASTGCN) [ASTGCN], diffusion convolutional recurrent neural network (DCRNN) [DCRNN], spatio-temporal graph convolutional network (STGCN) [STGCN], and multivariate time series forecasting with graph neural network (MTGNN) [MTGNN].
@@
SQ5 (Benchmark datasets). The datasets mentioned in the selected papers are closely related to the specific case of study. Even though some benchmark datasets are listed in the “Generic” thematic group, there is no common standard dataset for the entire spatio-temporal GNN research community. Hence, it would be beneficial for the community to agree on, and start adopting, some of the most commonly used datasets, with the goal of developing a shared benchmark dataset. A possible common benchmark could be the traffic one. Hopefully, this review could serve this purpose by providing comprehensive tables of the results of the models on different datasets to facilitate comparison between spatio-temporal GNN models.
@@
A meta-analysis of papers from different areas indicates that some thematic groups have seen more intense research on GNN models than others. For instance, the papers in the “Mobility” and the “Generic” groups present more complex models than others, this being likely due to the more extensive research conducted in these fields, as also evidenced by the larger number of published papers. In contrast, the “Finance” and “Health” thematic groups have less papers, and often have only a limited number of benchmarks available for comparison. Another notable example is the “Energy” group, which has a relatively large number of papers, particularly in 2024. However, the energy community tends to align more closely with traditional econometrics or non-GNN machine learning models, and the proposed GNN models in this field are generally less complex than those found in other groups. This suggests that some fields are currently more prone to the development of innovative and complex models, whereas others require a greater focus on adapting existing models to real-world problems. Furthermore, papers in the “Energy” group often test their models on proprietary dataset, which further complicates the development of models across the entire group. Consequently, we expect some form of cross-pollination between the communities, which is likely to reduce these differences over time.
A final aspect to highlight is the importance of reproducibility and comparability of these studies, which is essential for the advancement of scientific knowledge. In some fields, a greater attention to reproducibility can be observed, also thanks to the availability of datasets, a larger number of benchmarks, and more detailed descriptions of the models. Tab. 40 shows the average number of datasets and benchmark models used, as well as the percentage of papers reporting the source code for each group.

### v4

@@ 5.7.5 Results
The most common error metrics used in the evaluation of the models in the “Generic” group are, in order, the mean absolute error (MAE), correlation coefficient (CORR), root mean squared error (RMSE), root relative squared error (RRSE), mean absolute percentage error (MAPE), and mean squared error (MSE). Differently from other metrics, where lower values indicate higher accuracy, the correlation coefficient (CORR) quantifies the strength of the the linear relationship between predicted and actual values, with higher values indicating a better performance. This subsection about results compares the results obtained for the Electricity consumption, Exchange rate, Solar energy and Traffic datasets. The results of the proposed models and benchmarks on the “Mobility” datasets were already reported with a in Subsec. 5.5.5.
@@
DPGNN [240] 0.0720 0.9527 0.0810 0.9422 0.0893 0.9357 0.0935 0.9291
DSTGN [241] 0.0713 0.9518 0.0821 0.9424 0.0887 0.9357 - -
DSTIGNN [99] 0.0733 0.9515 0.0820 0.9412 0.0899 0.9338 0.0954 0.9289
ESG [448] 0.0718 0.9494 0.0844 0.9372 0.0898 0.9321 0.0962 0.9279
@@
MAGL [268] 0.0728 0.9476 0.0824 0.9377 0.0893 0.9299 0.0951 0.9240
MAGNN [40] 0.0745 0.9476 0.0876 0.9323 0.0908 0.9282 0.0963 0.9217
MDST-GNN [138] 0.0738 0.9454 0.0833 0.9346 0.0884 0.9264 0.0922 0.9222
MTGNN [424] 0.0745 0.9474 0.0878 0.9316 0.0916 0.9278 0.0953 0.9234
MTGODE [177] 0.0736 0.9430 0.0809 0.9340 0.0891 0.9279 - -
@@
The accuracy of the various models varies depending on the studied forecasting horizon but, overall, one can notice that the GNN models perform better than the others, with the static and dynamic graph learning network (SDGL) [243] and the temporal polynomial graph neural network (TPGNN) [271] being the most accurate. The SDGL model, proposed by Li et al. in 2023, first employs a static and dynamic graph learning module that incorporates graph regularization to control the smoothness, connectivity, and sparsity of the learned graph. The graph input then passes through gated temporal convolutional networks and graph convolution modules, which are equipped with residual and skip connections that deliver the information to the output layer. The TPGNN, proposed by Liu et al. in 2022, builds a series of matrix polynomials that share a common matrix basis. The graph learning module first captures the overall correlations using a static matrix basis. Then, a set of time-varying coefficients and the matrix basis are used to create a dynamic graph. Overall, the TPGNN model uses an encoder-decoder architecture and makes predictions in an auto-regressive manner.
@@
DPGNN [240] 0.0184 0.9797 0.0246 0.9715 0.0346 0.9564 0.0456 0.9386
DSTGN [241] 0.0179 0.9782 0.0250 0.9715 0.0346 0.9562 - -
DSTIGNN [99] 0.0173 0.9818 0.0244 0.9723 0.0333 0.9571 0.0430 0.9407
ESG [448] 0.0181 0.9792 0.0246 0.9717 0.0345 0.9564 0.0468 0.9392
@@
MAGL [268] 0.0178 0.9797 0.0246 0.9716 0.0337 0.9567 0.0444 0.9361
MAGNN [40] 0.0183 0.9778 0.0246 0.9712 0.0343 0.9557 0.0474 0.9339
MDST-GNN [138] 0.0172 0.9811 0.0245 0.9727 0.0337 0.9578 0.0431 0.9392
MTGNN [424] 0.0194 0.9786 0.0259 0.9708 0.0349 0.9551 0.0456 0.9372
@@
Also in this case, GNN models are the best. The so-called discrete graph structure learning for time series model (GTS), proposed in 2021 by Shang et al. in Ref. [338] for traffic and sensor network forecasting, and used as a benchmark in Ref. [243], demonstrates itself to have a superior performance. It includes a probabilistic graph learning module that optimizes the mean performance over the graph distribution, and a recurrent graph convolutional module to generate the forecasts.
@@
DPGNN [240] 0.1794 0.9850 0.2364 0.9724 0.3003 0.9543 0.4091 0.9117
DSTGN [241] 0.1787 0.9867 0.2358 0.9721 0.3079 0.9513 - -
DSTIGNN [99] 0.1684 0.9869 0.2165 0.9773 0.2863 0.9584 0.4064 0.9118
ESG [448] 0.1708 0.9865 0.2278 0.9743 0.3073 0.9519 0.4101 0.9100
@@
MAGL [268] 0.1678 0.9869 0.2220 0.9759 0.2998 0.9550 0.4188 0.9057
MAGNN [40] 0.1771 0.9853 0.2361 0.9724 0.3015 0.9539 0.4108 0.9097
MDST-GNN [138] 0.1764 0.9855 0.2321 0.9735 0.3082 0.9519 0.4119 0.9103
MTGNN [424] 0.1778 0.9852 0.2348 0.9726 0.3109 0.9509 0.4270 0.9031
MTGODE [177] 0.1693 0.9868 0.2171 0.9771 0.2901 0.9577 - -
@@
The results indicate that the dynamic spatiotemporal interactive graph neural network (DSTIGNN) model [99] is the most accurate on the solar energy dataset. The DSTIGNN model, introduced by Gao et al. in 2021, employs a hierarchical structure consisting of a spatio-temporal interactive learning module, a dynamic graph inference module, multiple spatio-temporal layers, and an output module. The spatio-temporal interactive learning module splits the time series into two subsequences to capture both temporal and spatial correlations. Then, the graph inference module learns a dynamic adjacency matrix for the graph. Once the graph structure is determined, a stack of graph convolutional layers, enhanced with residual connections, along with a convolutional output module, generate the final forecast.
@@
MAGL [268] 0.4058 0.9013 0.4189 0.8935 0.4308 0.8871 0.4417 0.8810
MAGNN [40] 0.4097 0.8992 0.4555 0.8753 0.4423 0.8815 0.4434 0.8813
MDST-GNN [138] 0.4162 0.8958 0.4461 0.8803 0.4377 0.8841 0.4452 0.8792
MTGNN [424] 0.4162 0.8963 0.4754 0.8667 0.4461 0.8794 0.4535 0.8810
MTGODE [177] 0.4127 0.9000 0.4259 0.8945 0.4329 0.8899 - -
@@
The accuracy of the models varies by horizon and metric. Overall, the best models are the temporal polynomial graph neural network (TPGNN) [271] already discussed for the Electricity consumption dataset, and the temporal decomposition enhanced graph neural network for multivariate time series forecasting (TDG4-MSF) [287]. The TDG4-MSF model, proposed by Miao et al. in 2021, is composed of four components: temporal decomposition enhanced representation learning for extracting short and long term periodic patterns, graph structure learning, gated graph neural network depending on linear transformation for representation learning, and MLP-based output layer.
There are also other pairs of papers whose results can be compared. However, due to space limitations, we do not include tables comparing the results of just two papers in this “Generic” subsection. Instead, the reader is referred to the studies Refs. [432] and [451] for the CSSE COVID-19 and ECG 5000 datasets; Refs. [50] and [51] for the Bytom dataset; Refs. [504] and [153] for the Jena weather dataset.
@@ 5.8 Other topics
This final subsection includes 34 papers that fall outside the previously discussed groups. The case studies include cellular traffic prediction, modulation classification, and other forecasting or classification problems. Given the heterogeneity of the papers, only a general overview is provided here, and no comparisons of the approaches or discussions of benchmarks are presented.
@@ 5.8.1 Overview
The interest in GNNs has grown so extensively that the selected themes are actually not sufficient to cover all relevant application domains. However, since only a few papers would belong to each potential additional theme, they are grouped under the label “Other topics”. A frequently discussed topic in this “Other topics” category is cellular traffic prediction [404, 34, 459, 170, 254, 108, 252], in terms of SMS messages, calls, internet connections, phone’s traffic statistics or locations, which are essential for the cellular network resource management system. Another topic included is modulation classification [380, 435, 6, 487, 454], which is the process of determining the modulation used at the transmitter based on observations of the received signal [124]. Four further areas of investigation to be added are the study of product demand and user preferences [92, 501] to enable personalized experiences and recommendations, anomalous IoT intrusion detection [3, 394], encrypted traffic classification [70, 475].
Other applications discussed in the papers of this subsection are urban spatio-temporal event prediction [175], prediction of the quality of operational processes [492], seismic intensity forecasting [16] (with source code available at https://github.com/StefanBloemheuvel/GCNTimeseriesRegression), prediction of spatio-temporal dynamics with complex structures [410], tailing dam monitoring [328], subsurface production forecasting [266], time series prediction for data centers maintenance [344], hydraulic runoff [440], water demand forecasting [465], crime prediction [519] (with source code available at https://github.com/ZJUDataIntelligence/HDM-GNN), polypropylene production data [130], tungsten flotation process [399] (with source code available at https://github.com/ywmcsu/SASGNN), talent and demand supply for labor market [121] (with source code available at https://github.com/gzn00417/DH-GEM), anomaly detection in microservice systems [150] (with source code available at https://github.com/ant-research/microservice_system_twin_graph_based_anomaly_detection ), and emergency supply forecasting for logistics systems [255].
@@ 5.8.2 Datasets
The two datasets for modulation classification used in Refs. [380, 435, 6, 454] are RML2016.10a (https://www.kaggle.com/datasets/raindrops12/rml201610a, source paper: http://dx.doi.org/10.1007/978-3-319-44188-7_16) and RML2016.10b (https://www.kaggle.com/datasets/marwanabudeeb/rml201610b, source paper: https://pubs.gnuradio.org/index.php/grcon/article/view/11). They include 220,000 and 1.2 million samples respectively, in many modulation types. Each modulation type has 20 levels of signal-to-noise ratios at 2 dB intervals from -20 dB to 18 dB.
As for the cellular traffic prediction dataset used in Refs. [303, 404, 34, 170, 108], it was proposed by Telecom Italia and MIT Media Lab during the “Telecom Italia Big Data Challenge”, and it is available at https://dataverse.harvard.edu/dataverse/bigdatachallenge. The investigated geographical area is divided in a grid, and each cell records the number of SMS messages, calls, and wireless network traffic data in a 10 minutes interval.
Finally, the Bot-IoT dataset used in Refs. [3, 394] and available at https://www.unsw.adfa.edu.au/unsw-canberra-cyber/cybersecurity/ADFA-NB15-Datasets/bot_iot.php, was created by designing a realistic network environment in the Cyber Range Lab of UNSW Canberra, and incorporates a combination of normal and botnet traffic.
@@ 6 Discussion
As stated in the Introduction, the aim of this SLR is to provide a comprehensive but also detailed overview of spatio-temporal GNN models for time series classification and forecasting used across different fields, assessing their performance and highlighting patterns that could guide future research. Existing literature was collected and synthesized to provide both an overview and a distillation of datasets, models, and tables of results, addressing the two sets of research questions posed in the Introduction.
Although our intention was to provide a broad and general perspective on widely used spatio-temporal GNN models and identify patterns that could be generalized across different domains and datasets, our analysis revealed that such generalizability is highly elusive. This characteristics stems from the primary strength of spatio-temporal GNNs: the possibility of explicitly modeling relationships between different variables, which is inherently dataset-dependent and highly heterogeneous across different applications. For instance, models for human mobility (e.g., traffic flow, epidemic spread) share intrinsic similarities but differ in temporal scales, spatial resolution, and interaction types. Similarly, environmental applications such as weather forecasting and energy demand prediction are both influenced by climatic factors but rely on different variables and relationships.
We believe that a comprehensive perspective can be drawn after such an extensive examination of the available literature. Hence, we attempt now to discuss common patterns and insights that are not immediately apparent across different domains. In this section, we first address the research questions posed in the Introduction, then summarize prevailing research trends, and finally present a meta-analysis aimed at uncovering common insights.
6.1 Research questions
@@
SQ2 (Graph construction). Most of the selected papers focus on pre-defined graph structures (when available), with the objective of extracting the maximum amount of information and enhancing the interpretability of the model. However, there is a recent growing interest in models that learn the graph structure and the edge weights themselves. While this trend toward adaptive graph learning is expected to gain momentum, the literature remains divided on whether learned or pre-defined adjacency matrices are preferable. The choice between these approaches should be based on the availability of domain knowledge and model performance requirements, requiring further research to establish clearer guidelines.
@@
SQ3 (Taxonomy). As observed in Ref. [467], there are two main approaches for the design of spatio-temporal GNNs: one that treats the spatial and temporal components in separate modules, and another that integrates and processes them together. The analysis of the collected papers reveals that the most common approach in literature is that of separate modules. Specifically, researchers very often address the spatial and temporal aspects of the problem independently from each other, and focus on each module separately. Hence, the proposed taxonomy of GNN models takes its main perspective from the point of view of the spatial component. The review indicates that convolutional approaches are the most prevalent, followed by attentional methods. Approximately 60% of the models are purely convolutional, 22% are purely attentional, and 10% are hybrid convolutional-attentional. Additional statistics related to taxonomy distributions by year and research field are available in the GitHub repository associated with this review. With regard to the temporal component, the recurrent structures of GRUs and the attention mechanism are widely used.
@@
SQ4 (Benchmark models). As for the benchmark models, there are many options available, and their choice depends on the specific application. In thematic groups such as “Energy”, “Finance”, “Health”, and “Predictive monitoring”, the focus tends to be on simpler statistical and non-GNN machine learning benchmarks, with only a limited use of GNN benchmarks. In the “Generic” group, many recent machine learning benchmark models based on the Transformer architecture have emerged. Notably, in “Mobility”, “Environment”, and “Generic” groups (which are also the most investigated fields), there are many reference GNN benchmark models. The most widely used GNN benchmarks, spanning all thematic groups, include: the graph convolutional neural network with long-short term memory (GC-LSTM) [316], spatio-temporal graph convolutional network (STGCN) [457], attention-based spatial-temporal graph convolutional network (ASTGCN) [116], diffusion convolutional recurrent neural network (DCRNN) [237], and multivariate time series forecasting with graph neural network (MTGNN) [424]. These GNN benchmarks, mostly published between 2018 and 2020, represent some of the earliest and most cited works in spatio-temporal GNN literature. Almost all of them employ graph convolutional mechanisms, often Chebyshev-based or similar variants, reflecting the state-of-the-art at the time. In most cases, the adjacency matrix is pre-defined, often based on distances between measurement sites, whereas just one approach incorporates a learned adjacency matrix (i.e., MTGNN, which is one of the first models to use self-learned graph structures). In terms of application domains, four out of five of these benchmarks were originally developed to focus on “Mobility” or “Generic” tasks, with only GC-LSTM originally addressing environmental applications. Among these benchmarks, ASTGCN also explores attention mechanisms alongside graph convolutions and constitutes one of the earliest hybrid architectures. Overall, these benchmarks illustrate the early development of GNNs: graph convolution dominates, adjacency matrices are usually static, and mobility-related tasks are the most commonly studied.
@@
SQ5 (Benchmark datasets). The datasets mentioned in the selected papers are closely related to the specific case of study. Even though some benchmark datasets are listed in the “Generic” thematic group, there is no common standard dataset for the entire spatio-temporal GNN research community. Hence, it would be beneficial for the community to agree on, and start adopting, some of the most commonly used datasets, with the goal of developing a shared benchmark dataset. Some possible common benchmark could be the ones presented in the “Mobility” and the “Generic” thematic group. Hopefully, this review could serve this purpose by providing comprehensive tables of the results of the models on different datasets to facilitate comparison between spatio-temporal GNN models.
@@
6.2 Observed trends
Our analysis of papers from different domains indicates that some thematic groups have seen more intense research on GNN models than others. For instance, the papers in the “Mobility” and “Generic” groups generally present more complex models, likely reflecting the greater research effort in these fields, as evidenced by the larger number of published papers. In contrast, the “Finance” and “Health” thematic groups have fewer papers and often offer only a limited number of benchmarks for comparison. Another notable example is the “Energy” group, which has a relatively large number of papers, particularly in 2024. However, the related community tends to rely more on traditional econometrics or non-GNN machine learning approaches, and the GNN models proposed are generally less complex than those in other domains. Moreover, many papers in the “Energy” thematic group test their models on proprietary datasets, which further complicates model development across the group. Overall, this analysis suggests that some fields are more prone to the development of innovative and complex models, whereas others focus on adapting existing models to specific real-world problems. Consequently, we expect some form of cross-pollination between the communities, which is likely to reduce these differences over time. For readers interested in a more detailed view of how GNN taxonomies evolve across fields and over time, interactive charts are available in the GitHub repository accompanying this paper.
Another important aspect related to research trends is the focus on reproducibility and comparability of studies, which, more in general, is essential for advancing scientific knowledge. Some fields pay greater attention to reproducibility, facilitated by the availability of datasets, a larger number of benchmarks, and more detailed model descriptions. Tab. 40 shows the average number of datasets and benchmark models used, as well as the percentage of papers reporting the source code for each group.